Operations Report – April 15th

Supplemental Operations Report 15-Apr-2023

Name of person filing report: Sergii Iakymov

Reason for Report: Routine.

Non-Nominal Systems: Robotic observatory.
Action taken for non-nominal systems: New parts for Robotic observatory are received and scheduled to be installed during Crew 261 rotation.

Power system:
Solar:SOC Last 24 hours: Max 100%; Min 70 %; Avg 82.4%.
VDC Last 24 hours: Max 59.15V; Min 44.06; Avg 51.65V.
On April 14th all three pairs of breakers in solar chargers input were tripped, for unknown reason. After reset all work is nominal.

Power System Notes: Generator run time about 11:30pm – 7am. On April 13th and 14th start time about 10-11pm due to clouds and crew final cleaning. Hours – 3695.8. Serviced on April 15th.

Propane Reading Station Tank: 56%
Propane Reading Director Tank: 73%
Propane Reading Intern Tank: 72%
Propane Reading Generator Tank: 55%

Water (Static Tank) – 200 gallons
Water in GreenHab – less than 40 gallons
Water (Outpost tank) – 353 gallons
Water in Science Dome: 0 gallons
Hab Toilet Tank emptied: n/a

Sojourner rover used: No
Hours: 190.4
Beginning Charge: 100
Ending Charge: 100
Currently Charging: Yes
Notes on Rovers: All are plugged in, charged. All corroded battery wires were received and replaced. Sojourner needs one additional short battery wire to be replaced. It will be ordered as soon as possible.
ATV’s Used: (Honda, 350.1, 350.2, 300): None
Reason for use: n/a
Oil Added? No
ATV Fuel Used: 0 Gals
Ethanol Free Gasoline — 0 Gallons
# Hours the ATVs were Used today: 0
Notes on ATV: All parked at the outpost and awaiting to be taken to service.

HabCar used and why, where? To Loa for repair. Malfunctioned alternator was replaced and Hab Car is scheduled to be picked up next Wednesday, April 19th.

Crew Car used and why, where? Yes. To Hanksville for water runs and supplies, by crew to Grand Junction.

General notes and comments: nothing to report

Dual Split in Science Dome: off due to warmer weather.

Summary of Internet: Nominal

Summary of suits and radios: Nominal

EVA COMMS: Nominal

Campus wide inspection, if action taken, what and why: nothing to report

Summary of General Operations: All nominal.

Summary of Hab operations: Nominal. The heater air filter was replaced today. One mouse was caught on the lower deck. First AID kit was resupplied with Dramamine.

Summary of GreenHab Operations: Nominal.

Summary of SciDome Operations: Nominal.

Summary of any Observatory Issues: New parts for Robotic observatory are received and scheduled to be installed during Crew 261 rotation .

Summary of RAM Operations: Nothing to report

Summary of Outpost Operations:

Director’s trailer:
1) Toilet flush is leaking from time to time and will need to be replaced. Parts are on hand.
Interns trailer:
1) Bathroom faucet has been replaced.
2) Trailer cleaned and heater set up to 55 F due to warm weather.

Summary of Health and Safety Issues: Nominal.

Questions, Concerns, Supplies needed and Requests: Nothing to report

Copyright © The Mars Society. All rights reserved. | Main Site